Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, December 14, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 14, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/14/98 (Monday)

Segment 1 report for 12/14/98: The equipment at Emidio Pump Station was turned on and tested. On the west side of the Stream Crossing #4 bore, pipe that was tied-in from EMP 8.98 to 8.94 was coated, on the east side pipe was being welded from EMP 9.04 to 9.08 and a pullbox was installed at EMP 8.98. The shoring boxes were removed on the west side of the Tecuya Creek/I-5 bore (EMP 9.67) and backfill began. Pipe was being welded, x-rayed and coated at the California Aqueduct bore (EMP 12.28). The bore operations continued at the Stream Crossing #9 bore (15.35) and the Tejon Bench/I-5 bore (EMP 16.49). Electrical work continued at Grapevine Pump Station. On Grapevine Rd., shading continued from EMP 16.07 to 16.01. Due to intermittent heavy rain no work was done on Tejon Bench, at the Warbler skip on Lebec Rd., and at the Cuddy Creek Tributary bore (EMP 25.47)

Segment 2B (Castaic to Sylmar): Report for 12-14-98. Fibber optic cable was pulled between Magic Mtn Parkway and Valencia Blvd. In addition, splicing crews worked connecting cable at EMP 69.7. At Weldon Cyn, crews continued removing wood shoring and jacks from the right of way. In addition, asphalt and spoils were hauled to a local landfill. Near EMP 72.4, crews worked to complete the electrical tie in to the vault. Both trenching and backfill were accomplished with paving planned for Tuesday. Further paving was completed at several locations south of Sierra Hwy while steel plates were removed from the right of way at Sunshine Cyn Rd. In addition, a sweeper was active along the right of way near Weldon Cyn to vacuum gutters clean of asphalt debris.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 14 December 1998: 3A: Paving crews worked on final paving in Glendale today. Backfill and FOC crews continued to work at San Fernando and Hollywood Way Bridges. Ditch crews dug out the connection between the pipe and the bore at Tuxford Street. Welding, x-ray, and coating crews worked on the pipe to complete the skip at Tuxford Street. Coating crews also worked on placing insulation pieces on the various bridge crossings. 3C: Welding, x-ray, and coating crews completed their work on the tie in to the bore and skip at Hubbard Avenue. Shading and FOC crews began to work on the Hubbard area. Fiber optic crews worked on the north end pulling cable and digging assist pits. Backfill crews continued clean up work. Welders continued work for the pipe needed to fill the line for the hydro test.

Inspection Summar

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page